two squares are chosen at random on a chess board findthe probability that they have common side?


 


ans:i got the ans as 38/3451..........is this correct ? if not how to do?


 

Grade:

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

 

Dear anesh krishna,

 64 no of squares and we can choose first square by 64 ways

and second by 63ways.

There are a total of 64 * 63 = 4032



corners have only 2 adjacent squares. we have 4 corners.

total we can choose using corner square are=4*2

another 24 squares along the boundaries they have 3 adjacent squares.

total no= 24*3

the remaing squares are 34 they have 4 adjacent sides.

total= 36*4

number of combinations = (4 * 2) + (24 * 3) + (36 * 4) = 224

 = .

 

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar Singh

IIT Bombay